OpenStudy (anonymous):

A farmer wants to fence an area of 6 million square feet in a rectangular field and then divide it in half with a fence parallel to one of the sides of the rectangle. Find the dimensions of the rectangle he should use to minimize the cost of fencing. List the dimensions in non-decreasing order.

OpenStudy (anonymous):

would it be 3,000 and 2,000?

OpenStudy (anonymous):

I might have inputed the answers differently. I'll try again, thank you!

OpenStudy (anonymous):

If you want me to double check your work out solution, I'll be happy, but don't ask me for the answer, I myself haven't computed it yet!

OpenStudy (anonymous):

Here's what I did: 2p + 3q LW = 6 x 10^6 sqft. =>2p + [3(6x10^6) =2p + (18 x 10^6) (p^-1) dP/dp = 2+ (18 x 10^6)(-1)(p^-1) => 2 + [(18 x 10^6)/(p^2)] = 0 2p^2 - 18 x 10^6 = 0 p^2 = sqrt 36 and sqrt 10^6 = 6 x 1,000 THUS, p= 6,000ft q = (6 x 10^6)/ 6,000 = 1,000 ft Not sure if this is correct.

OpenStudy (anonymous):

I got w = 2,000 ft and L = 3,000 ft

OpenStudy (anonymous):

the other way was: 3x +2y A =xy = 6,000,000 y = 6,000, 000/x L = 2x + 12,000,000/x 3- (12,000,000/x^2) = 0 3 = 12, 000,000/x^2 sqrt x^2 = sqrt 4,000,000 x = 2,000 and y = 3,000 What's non-decreasing order? 2000ft, 3000ft?

OpenStudy (anonymous):

so i guess that was the correct way. AWW, cute!

OpenStudy (anonymous):

Okay, I got it. Thank you so much!

OpenStudy (anonymous):

your way is the best way and probably the easiest. thank you! :)

OpenStudy (anonymous):

Now I'm happy! You're the first one compliment my way is the easiest one ;)

OpenStudy (anonymous):

:) It helped me get the correct answer, so I'm pretty excited. I was on my last try as well!